OpenStudy (moonlitfate):

More optimization problems! A rectangular page is to contain 100 square inches of print. The margins on each side are 1 inch. Find the dimensions of the page such that the least amount of paper is used.

OpenStudy (unklerhaukus):

start with some equations

OpenStudy (moonlitfate):

Okay, well, I know that the area is going to be: A=x*y ; and I think the page looks like this: 

OpenStudy (moonlitfate):

Am, I kind of right? :x

OpenStudy (moonlitfate):

These problems are kinda confusing.

OpenStudy (unklerhaukus):

if the dimensions of the page are x, y. then the text will have area (x-2)(y-2)=100

OpenStudy (moonlitfate):

Ahh, the -2 is just because 2 inches are taken off each side for margins, got it.

OpenStudy (moonlitfate):

So, I think next I would have to solve for one of the variables?

OpenStudy (unklerhaukus):

yess...

OpenStudy (moonlitfate):

A=x*y Constraint --> A=x*y=100; So y = 100-x ? :D

OpenStudy (unklerhaukus):

the area of the print was (x-2)*(y-2)=100 [solve this for y] the area of the page was A=x*y

OpenStudy (moonlitfate):

Hmm... xy-2x-2y+4 = 100; I have no idea how my mind is becoming flustered by such a simple thing, ugh. =_=

OpenStudy (agent0smith):

@MoonlitFate don't foil it. We can solve for y much easier: \[\Large (x-2)*(y-2)=100\] what happens if we divide both sides by (x-2)

OpenStudy (moonlitfate):

Wow, I fail. Lol.\[y-2 = \frac{ 100 }{ x-2 }\]

OpenStudy (moonlitfate):

+2 no -2 on the right.

OpenStudy (agent0smith):

Correct, now add 2 to both sides (or is that what you meant with your last post..?)

OpenStudy (moonlitfate):

\[y = \frac{ 100 }{ x+2 }+2\]

OpenStudy (unklerhaukus):

how did that two in the denominator change sign?

OpenStudy (moonlitfate):

Apparently my brain is turning to mush. \[y = \frac{ 100 }{ x-2 }+2\] Am I good now?

OpenStudy (unklerhaukus):

that is good, now plug this into A=x*y

OpenStudy (moonlitfate):

All right.\[A=x(\frac{ 100 }{ x-2 }+2)\]\[=\frac{ 100x }{ x-2 }+2x\]

OpenStudy (unklerhaukus):

ok, good, can you simplify that fraction a little?

OpenStudy (moonlitfate):

\[-50+2x?\]

OpenStudy (unklerhaukus):

um , you cant do that/

OpenStudy (unklerhaukus):

maybe just leave the fraction as it was before, can you find the derivative of A

OpenStudy (moonlitfate):

Oh, oops. D: Sorry. Yeah, I can find the derivative.

OpenStudy (moonlitfate):

\[A'(x) = 2-\frac{ 200 }{ (x-2)^2 }\]

OpenStudy (unklerhaukus):

good

OpenStudy (agent0smith):

Now, since we want to minimize A, you need to find where A'(x) = 0, so solve for x \[\Large A'(x) = 2-\frac{ 200 }{ (x-2)^2 } = 0\] We might have to check the sign of the second derivative to confirm it's a minimum, but for now, just solve for x.

OpenStudy (moonlitfate):

All right. x=12 or x = -8

OpenStudy (unklerhaukus):

Great stuff, now only one of those values for x is going to the length of a rectangle, right?

OpenStudy (unklerhaukus):

plug this back into (x-2)*(y-2)=100 and solve for y and your done .

OpenStudy (moonlitfate):

Well, you can't have a negative value for height. :p

OpenStudy (unklerhaukus):

thats right

OpenStudy (moonlitfate):

I got that y would also be equal to 12. :)

OpenStudy (unklerhaukus):

\[\large\color{red}{\checkmark\checkmark}\]